OpenStudy (krazzikitten69):

1. Graph the function. Then identify the domain and range of the function. y= -1/2x^2 2. A person walking across a bridge accidentally drops an orange into the river below from a height of 40 ft. The function h= -16t^2+40 give the orange's approximate height h above the water, in feet, after t seconds. Graph the function. In how many seconds will the orange hit the water? Ps. Explain both answers

OpenStudy (johnweldon1993):

 That is the graph...it is slightly "fatter" than just -x^2 Domain....what numbers on the 'x' axis will this reach? (ps..the graph doesn't stop there...those go on forever)

OpenStudy (krazzikitten69):

I don't know:|

OpenStudy (johnweldon1993):

That's alright :) Alright so...it will reach ALL of the numbers on the x-axis....why? well because like I said that parabola will go on forever...always going always going...of course going DOWN...but still...so it will reach every number on that x-axis eventually.. so the domain is (-infinity , infinity)

OpenStudy (johnweldon1993):

Also you can think about it like this.....is there any number you can plug into -1/2x^2 that will make it Undefined? that will make it not a real number? no...so the domain would be any number (-infinity , infinity)

OpenStudy (johnweldon1993):

Now...RANGE... Range is the y-axis.... See how the graph starts from the origin (0,0) and just goes down? Well (sadly) it never touches any of the positive y-axis...it starts at 0...and goes down to -infinity... So the range...is (0 , -infinity)

OpenStudy (johnweldon1993):

Another way to think about THAT one...is ANY number you plug into that equation -1/2x^2 lets do 10 -1/2(10)^2 = -1/2(100) = -50 What happened there? it became a negative number... even if we plug in -10 -1/2(-10)^2 = -1/2(100) = -50 Again -50...why? because when you square a negative number...it becomes positive...but we still had to deal with that -1/2 part in front... So see...this equation will never be a POSITIVE number...so its range again is (0 , -infinity)

OpenStudy (johnweldon1993):

Let me know if that makes sense @krazzikitten69 :)

OpenStudy (krazzikitten69):

Kind of.. so what would I say my answer would be for a quick explanation

OpenStudy (johnweldon1993):

You would say Domain = (-infinity , infinity) And Range = (0 , -infinity)

OpenStudy (krazzikitten69):

ohh okay its making sense now:) Can you help me with the other?

OpenStudy (johnweldon1993):

Sure :) so the graph would look something like